How many ounces of pure water must be added to 80 oz of a 9% salt solution to make a 6% salt solution?

A.
120 oz

B.
60 oz

C.
50 oz

D.
40 oz



Answer by stanbon(48551)

How many ounces of pure water must be added to 80 oz of a 9% salt solution to make a 6% salt solution?
-----------
9% solution DATA:
Amt. = 80 oz ; salt = 0.09*80 = 7.2 oz
--------------------
Water DATA:
Amt. = x oz ; salt = 0
-------------
Mixture DATA:
Amt. = (80+x)oz ; salt = 0.06(80+x) = 4.8 + 0.06x
-----------------
EQUATION:
salt + salt = salt
7.2 + 0 = 4.8+0.06x
0.06x = 2.4
x = 40 oz (amt. of pure water that must be added)
===============
Cheers,
Stan H.